OpenStudy (yourboydusty):

The circle below has a radius of 10 cm. What is the area of the shaded region? If necessary, round your answer to two decimal places. Do not include units in your answer.

OpenStudy (yourboydusty):

its 56 right?

OpenStudy (skullpatrol):

how did you get that?

OpenStudy (yourboydusty):

i don't even know lol i think i subtracted and divided?

OpenStudy (yourboydusty):

i lost my paper work lol

OpenStudy (skullpatrol):

start with the formula for the area of a circle...

OpenStudy (yourboydusty):

a= pi,r,2

OpenStudy (skullpatrol):

ok, calculate the entire area

OpenStudy (yourboydusty):

10*10=100*pi=314.159265359

OpenStudy (skullpatrol):

good, now how many degrees are in a circle?

OpenStudy (yourboydusty):

360?

OpenStudy (skullpatrol):

yes

OpenStudy (skullpatrol):

in this question you are given 120 degrees, right?

OpenStudy (yourboydusty):

yes

OpenStudy (skullpatrol):

how many times does 120 go into 360?

OpenStudy (yourboydusty):

3 times

OpenStudy (skullpatrol):

so take a third of the entire area, right?

OpenStudy (yourboydusty):

um yeah...

OpenStudy (skullpatrol):

if they gave you 180 degrees you would take 1/2 the area

OpenStudy (yourboydusty):

um not to be rude or anything but i really need answers....i have like 50 something questions to do and only a short amount of time....again srry really not trying to be rude :(

OpenStudy (skullpatrol):

10*10=100*pi=314.15926535 $$\dfrac{314.159265359}{3}$$

OpenStudy (skullpatrol):

that^ is your answer :-)

OpenStudy (yourboydusty):

104.72

OpenStudy (skullpatrol):

Good job!

OpenStudy (yourboydusty):

thank you!

OpenStudy (yourboydusty):

for your help

OpenStudy (skullpatrol):

Thanks for trying to learn :D